OpenStudy (maddie82502):

A bake sale has 12 cookies, 9 pies, 8 cakes, 4 tarts, and 3 cobblers to sell. What is the ratio of the total number of baked goods to cookies in simplest form?

OpenStudy (peach3113):

12+9+8+4+3=?

OpenStudy (maddie82502):

36

OpenStudy (peach3113):

Yes!

OpenStudy (peach3113):

So, 36/12

OpenStudy (peach3113):

But reduce.

OpenStudy (peach3113):

36/12 can be divided by 12 36÷12=3 12÷12=1 3/1 :D

OpenStudy (peach3113):

3 to 1, 3/1, or 3:1

OpenStudy (texaschic101):

nice explanation :)

OpenStudy (peach3113):

Thank you. :D

OpenStudy (maddie82502):

thank u @peach3113 I appreciate it

OpenStudy (peach3113):

@maddie82502 No problemo! :D I love ratios. If you want, you can tag me whenever you have another question. :P

OpenStudy (maddie82502):

mmk